OpenStudy (anonymous):

Two circles share the same center. The radius of the larger circle is “r” and the radius of the smaller circle is 5 units less than the radius of the larger circle. How can I find the area of the non-overlapping space? --- I am not sure of what the equation I would use is.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\pi r^2\] for the big one \[\pi(r-5)^2\] for the small and therefore \[\pi r^2-\pi(r-5)^2\] for the non overlap

OpenStudy (anonymous):

ok, thank you!